OpenStudy (anonymous):

How do I solve problems like this? "What is the denominator of the sixth term?" 1, 1/8, 1/27, 1/64, ... 1/n^3

OpenStudy (anonymous):

\[1, 1/8, 1/27, 1/64, ... 1/n^3\]

OpenStudy (anonymous):

@Directrix

OpenStudy (freckles):

You have on the bottoms this: 1^3,2^3,3^3,...

OpenStudy (freckles):

So the bottom for the sixth term would be?

OpenStudy (anonymous):

@freckles Sorry, I'm not sure I know what you mean by that.

OpenStudy (freckles):

Denominators is what I was calling bottoms

OpenStudy (anonymous):

Oh ok. I still am not sure what you mean. I haven't done any kind of math in a long time and I am just trying to refresh my memory. So I honestly have no idea how to do problems like this.

OpenStudy (freckles):

First term=1/1^3 Second term=1/2^3 Third term=1/3^3

OpenStudy (freckles):

The only thing that is changing is the base in the denominators And it keeps going up by one each time

OpenStudy (freckles):

Let me write it another way 1term=1/1^3 2term=1/2^3 3term=1/3^3 So 6term=?

OpenStudy (freckles):

Another example the 7th term=1/7^3

OpenStudy (freckles):

The 30th term would be 1/30^3 and the denominator of the 30th term would be 30^3